A quality control engineer at a potato chip company tests the bag filling machine by weighing bags of potato chips. Not every bag contains exactly the same weight. But if more than 18% of bags are over-filled then they stop production to fix the machine. They define over-filled to be more than 1 ounce above the weight on the package.

The engineer weighs 206 bags and finds that 50 of them are over-filled. He plans to test the hypotheses H0: p = 0.18 versus Ha: p > 0.18. What is the test statistic?

z =

(Please round your answer to two decimal places if necessary.)
Answer #1

Solution :

\hat p = x / n = 50 / 206 = 0.2427

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.2427 - 0.18 / [\sqrt(0.18 * 0.82) / 206]

= 2.34

z = 2.34